OpenStudy (anonymous):

Name the algebraic property demonstrated in the example below. If 4 = x + y, then x + y = 4. Reflexive Property Associative Property of Addition Symmetric Property Commutative Property of Multiplication

OpenStudy (anonymous):

Well just wach, ITS Symmetric Property!! Thanks to google........

OpenStudy (anonymous):

Its associative property of addition :/ I guess so :/

OpenStudy (anonymous):

Associative property of addition :)

OpenStudy (mathstudent55):

Here are the facts: Reflexive: a = a Associative: a + (b + c) = (a + b) + c Symmetric: If a = b, then b = a Commutative: a + b = b + a Now you decide.

OpenStudy (anonymous):

Sorry confused :/

OpenStudy (mathstudent55):

Hint: It's not associative.

OpenStudy (anonymous):

@mathstudent55 I guess @MATHAMATICAL is correct xD

OpenStudy (mathstudent55):

That leaves only 3 choices.

OpenStudy (anonymous):

It's symmetric

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

Hahaha cool :D

OpenStudy (anonymous):

Thanks for the medal xD